OpenStudy (anonymous):

Describe the transformations required to obtain the graph of the function f(x) from the graph of the function g(x). (x) = cos x/3; g(x) = cos x @agent0smith

OpenStudy (agent0smith):

like this?\[\Large f(x) = \cos (\frac{ x }{ 3 })\]

OpenStudy (anonymous):

i guess?

OpenStudy (agent0smith):

Well because this is different.\[\Large f(x) = \frac{ \cos x }{ 3 }\]

OpenStudy (anonymous):

The question is exactly as i wrote it

OpenStudy (anonymous):

so i honestly have no idea

OpenStudy (anonymous):

I think i got it thanks c: